If a force of 50N moves a load of 20kg up an inclined plane . Determine the magnitude of the force if the plane is inclined at an angle of 30 degree.
1
Expert's answer
2019-11-27T10:14:52-0500

The net force is


"F_{net }=mg\sin{30}-F"

"F_{net }=(20)(9.8)\sin{30}-50=48\ N"